A certain drug can be used to reduce the acid produced by the body and heal damage to the esophagus due to acid reflux. The manufacturer of the drug claims that more than

9393​%

of patients taking the drug are healed within 8 weeks. In clinical​ trials,

214214

of

229229

patients suffering from acid reflux disease were healed after 8 weeks. Test the​ manufacturer's claim at the

alpha equals 0.01α=0.01

level of significance.Click here to view the standard normal distribution table (page 1).

LOADING...

     Click here to view the standard normal distribution table (page 2).

LOADING...

Because np 0 left parenthesis 1 minus p 0 right parenthesisnp01−p0equals=nothing

not equals≠

less than<

greater than>

equals=

​10, the sample size is

less thanless than

greater thangreater than

​5% of the population​ size, and the sample

is given to not be random,

cannot be reasonably assumed to be random,

is given to be random,

can be reasonably assumed to be random,

the requirements for testing the hypothesis

are not

are

satisfied.

​(Round to one decimal place as​ needed.)
Answer #1

H0 : p = 0.93
Ha: p > 0.93

test statistics:
phat = 214/229 = 0.9345

z = (phat - p)/sqrt(p*(1-p)/n)
= ( 0.9345 - 0.93) /sqrt(0.93*(1-0.93)/229)
= 0.27

p value = 0.3948


Do not reject H0 as p value > 0.01